Ask your own question, for FREE!
Mathematics 11 Online
OpenStudy (anonymous):

When Ayla improves her game of darts the chances that a she hits a bullseye is 0.65. Assume that each throw is independent. a) What are the chances that three darts fired in succession will all hit bullseye? b) What is the probability that none will hit? c) What is the probability that at least one will hit? d) What is the probability that 2 will hit?

OpenStudy (anonymous):

assuming independence, multiply \[(.65)^3\] for the first one

OpenStudy (anonymous):

Can you put it in % form please?

OpenStudy (anonymous):

probability she doesn't hit a bull's eye is \(1-.65=.35\) so missing 3 times in a row is \((.35)^3\)

OpenStudy (anonymous):

i did not compute these numbers, i only wrote what you need to compute. once you compute them you can convert them to decimals by moving the decimal point two spaces to the right

OpenStudy (anonymous):

ok Can you help me with c and d?

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!